# RouteLLM: Learning to Route LLMs with Preference Data

### v3

@@ 4.2 Quantifying dataset and benchmark similarity
We attribute the difference in the performance of routers trained on the same dataset across different benchmarks to the differing distributions of evaluation data and training data. For each benchmark-dataset pair, we compute a benchmark-dataset similarity score in Table 4 indicating how well-represented evaluation data is in the training data, described in detail in Appendix C.
@@
A higher benchmark-dataset similarity score is correlated with stronger performance on that benchmark for routers trained using the corresponding dataset, as shown in Section 4.1. Dataset augmentation, be it using golden-labeled datasets or LLM-judge-labeled datasets, shifts the overall distribution of the preference data to be more in line with the benchmarks and increases the benchmark-dataset similarity score, which translates into performance improvements. This similarity score is also useful for understanding the relative performance of routers across different benchmarks: on the Arena dataset, the similarity score between MT bench and all datasets is noticeably greater than other benchmarks, which we believe explains the relatively stronger router performance on MT Bench as compared to GSM8K and MMLU. Benchmark-dataset similarity scores are a promising direction to systematically improve router performance in real-world use cases given knowledge about the query distribution.
4.3 Generalizing to other model pairs
We pick gpt-4-1106-preview [23] and Mixtral 8x7B [19] as representative strong and weak models for the above experiments. However, to demonstrate the generalizability of our framework to different model pairs, we report in this section our router performance on MT Bench when routed between Claude 3 Opus [6] and Llama 3 8B [5]. Importantly, we use the same routers without any retraining, and only replace the strong model and weak model routed to. These two models are also not present in our training data.
Train Set Method $CPT(50\%)$ $CPT(80\%)$ $APGR$
Random (95% CI) 47.23($\pm$5)77.08($\pm$5)0.494($\pm$0.03) (+0%)
$\mathcal{D}_{\text{arena}}$ BERT 60.61% 88.39% 0.475 (-3.9%)
Causal LLM 31.96% 59.83% 0.645 (+30.5%)
Matrix Factorization 24.84% 85.73% 0.605 (+22.5%)
SW Ranking 33.54% 80.31% 0.553 (+12%)
$\mathcal{D}_{\text{arena}}+\mathcal{D}_{\text{judge}}$ BERT 36.28% 50.83% 0.618 (+25.2%)
Causal LLM 40.46% 55.83% 0.625 (+26.5%)
Matrix Factorization 30.48% 41.81% 0.703 (+42.2%)
SW Ranking 31.67% 48.39% 0.667 (+35%)
Table 5: MT Bench results for Claude 3 Opus / Llama 3 8B. Our routers generalize very well across different strong and weak model pairs without any retraining.
Again, we observe strong results across all existing routers on MT Bench even when the model pair is replaced. Performance across all routers is comparable to with the original model pair. Results for both the new model pair and original model pair are still significantly stronger than random, with our routers requiring up to 30% less GPT-4 calls than random to achieve CPT(80%). These results suggest that our routers have learned some common characteristics of problems that can distinguish between strong and weak models, which generalize to new strong and weak model pairs without additional training.
@@ 4.4 Cost analysis
CPT 50% CPT 80%
MT Bench 3.66x (at 95% GPT-4 quality) 2.49x
MMLU 1.41x (at 92% GPT-4 quality) 1.14x
GSM8K 1.49x (at 87% GPT-4 quality) 1.27x
@@
We estimate the average cost of using GPT-4 and Mixtral 8x7B to be \$24.7 per million tokens and \$0.24 per million tokens respectively (see details in Appendix D). In Table 6, we show results of quantifying the cost savings achieved by our approach. We calculate the inverse of the ratio of GPT-4 calls utilized by our top-performing router relative to the random baseline because the cost of GPT-4 is the dominant factor in our analysis. Our routers achieve optimal cost savings of up to 3.66x, demonstrating that routing can significantly reduce cost while maintaining response quality.
@@ 4.5 Routing Overhead
SW Ranking \$37.36 2.9 \$0.39
Matrix Factorization \$1.42 155.16 \$0.8
@@
A concern with LLM routing is the overhead of routing as compared to using a single model. Therefore, we measure and report the overhead of our routers in Table 7 to demonstrate their practicality using randomly-sampled conversations from Chatbot Arena. For routers that requires GPUs, namely matrix factorization and the classifier methods, we utilize Google Cloud’s g2-standard-4 VM containing a single NVIDIA L4 GPU. For similarity-weighted ranking, we use Google Cloud’s CPU-only n2-standard-8 VM. Our GPU-based routers are currently much more efficient that our CPU-based routers, but we note that there is still much room for improvement in optimizing the throughput of our routers. However, even SW ranking, our most expensive router, introduces an additional cost of no more than 0.4% when compared to GPT-4 generation, as detailed in Appendix D.
@@ 5 Conclusion
We demonstrate strong routing performance by our routers across a variety of benchmarks, spanning open-ended question answering, humanities, and math problems. By intelligently routing queries between a strong model and weak model, our routers are able to achieve significant cost savings while maintaining a high response quality.
Our results also highlight the effectiveness of dataset augmentation in improving router performance. While training routers solely on the Arena dataset results in poor performance with MMLU and GSM8K, augmenting the training data with an LLM judge or in-domain data enables our routers to outperform the random baseline across all benchmarks. The largest performance gains occur when the training data closely resembles the evaluation data, as indicated by the benchmark-dataset similarity score. We believe that this framework provides a clear and scalable path to enhancing routing performance for specific use cases.
While our work demonstrates strong results, there are a few limitations. First, although we evaluate on a diverse set of benchmarks, real-world applications may have distributions that differ substantially from these benchmarks. To this end, we show that users can collect a small amount of in-domain data to improve performance for their specific use cases via dataset augmentation. Next, while we focus on the two-model routing setting in this work, a promising future direction would be to extend this work to multiple models. Finally, in our experiments, we observe that performance between different routers trained on the same dataset can vary widely on the same benchmark without a clear explanation—we leave further investigation into this for a future work.
@@ Appendix E Independent Benchmarks
In Figure 2, we present the performance of our best-performing routers on MT Bench as compared to Unify AI [4] and Martian [22], two existing commercial offerings for LLM routing. Here, we route between gpt-4-turbo-2024-04-09 [23] as our strong model, and mixtral-8x7b-instruct-v0.1 [19] or llama-2-70b-chat [29] as our weak model depending on the available models. For Unify AI, we select the best-performing router configuration on the user dashboard and use it for benchmarking. For Martian, we optimize for performance and specify the maximum cost per million tokens as \$10.45, approximating this value using public inference costs [24, 27] based on a 1:1 input:output token ratio such that 50% of calls are routed to GPT-4. Both the matrix factorization router and causal LLM routers perform very competitively when trained on $\mathcal{D}_{\text{arena}}$ augmented with $\mathcal{D}_{\text{judge}}$, achieving the same performance as these routers with up to 40% fewer calls routed to GPT-4.
@@ Appendix F Additional Plots
We include additional plots for all results presented in Section 4.1.
@@
Figure 6: MT Bench performance for all routers when routed to Claude 3 Opus and Llama 3 8B instead, without any retraining.

### v4

@@ 5.3 Quantifying dataset and benchmark similarity
We attribute the difference in the performance of routers trained on the same dataset across different benchmarks to the differing distributions of evaluation data and training data. For each benchmark-dataset pair, we compute a benchmark-dataset similarity score in Table 5 indicating how well-represented evaluation data is in the training data, detailed in Appendix C.
@@
A higher benchmark-dataset similarity score is correlated with stronger performance on that benchmark for routers trained using the corresponding dataset, as shown in Section 5.1. Dataset augmentation, be it using golden-labeled or LLM-judge-labeled datasets, shifts the overall distribution of the preference data to be more in line with the benchmarks and increases the benchmark-dataset similarity score, which translates into improved performance. This similarity score is also useful for understanding the relative performance of routers across different benchmarks: on $\mathcal{D}_{\text{arena}}$, the similarity score between MT Bench and all datasets is noticeably greater than other benchmarks, which we believe explains the relatively stronger router performance on MT Bench as compared to GSM8K and MMLU. Benchmark-dataset similarity scores are a promising direction for systematically improving router performance in real-world use cases, given knowledge about the query distribution.
@@ 5.4 Cost analysis
$CPT(50\%)$ $CPT(80\%)$
MT Bench 3.66 (95% GPT-4 quality) 2.49
MMLU 1.41 (92% GPT-4 quality) 1.14
GSM8K 1.49 (87% GPT-4 quality) 1.27
@@
We estimate the average cost of using GPT-4 and Mixtral 8x7B to be \$24.7 per million tokens and \$0.24 per million tokens respectively (detailed in Appendix D). Based on this, in Table 6, we quantify the cost savings achieved by our approach. To do so, we calculate the inverse of the ratio of GPT-4 calls made by our top-performing router relative to the random baseline because the cost of GPT-4 is the dominant factor in our analysis. The results show that our routers achieve cost savings of up to 3.66x, demonstrating that routing can significantly reduce cost while maintaining response quality.
@@ 5.5 Routing Overhead
SW Ranking \$39.26 2.9 \$0.39
Matrix Factorization \$3.32 155.16 \$0.8
@@
A concern with LLM routing is the overhead of routing as compared to using a single model. Therefore, we measure and report the overhead of our routers in Table 7 using randomly-sampled conversations from Chatbot Arena. For each router, we first profile the rate at which it can process requests, then use the VM’s hourly cost to calculate the cost overhead. For routers that require embeddings, we also include the cost of embedding generation based on the average input length of the training set. For routers that use GPUs, namely matrix factorization and the classifier methods, we utilize Google Cloud’s g2-standard-4 VM containing a single NVIDIA L4 GPU. For similarity-weighted ranking, we use Google Cloud’s CPU-only n2-standard-8 VM.
Our GPU-based routers are currently much more efficient that our CPU-based routers, but we note that there is still much room for improvement in optimizing these routers. Based on the results, our most expensive router, SW ranking, currently adds an extra cost of no more than 0.4% as compared to GPT-4 generation (detailed in Appendix D), demonstrating the cost-effectiveness of these routers.
@@ 6 Conclusion
We demonstrate strong performance by our routers across a variety of benchmarks from open-ended question answering to humanities and math problems. By intelligently routing queries between a strong and weak model, our routers achieve significant cost savings and high response quality without excessive cost or latency overhead. We also show that our routers maintain their performance across multiple strong / weak model pairs without retraining–an important capability that if absent, would greatly limit usefulness.
Our results highlight the effectiveness of dataset augmentation in improving router performance. While training routers solely on $\mathcal{D}_{\text{arena}}$ results in poor performance on MMLU and GSM8K, augmenting the training data with an LLM judge or in-domain data enables our routers to outperform the random baseline across all benchmarks. The greatest performance gains occur when the training data closely resembles the evaluation data, as indicated by the benchmark-dataset similarity score. We believe that this framework provides a clear path towards improving routing performance for specific use cases.
While our work demonstrates strong results, there are a few limitations. First, although we evaluate on a diverse set of benchmarks, real-world applications may have distributions that differ substantially from these benchmarks. To this end, we show that users can collect a small amount of in-domain data to improve performance for their specific use cases via dataset augmentation. Next, while we focus on the two-model routing setting in this work, a promising future direction would be to extend this approach to multiple models. Finally, rather than there being a single best router for all queries, the decision of which router to use should be based holistically on latency and cost requirements, as well as the types of queries handled. In our experiments, we observe that performance between different routers trained on the same dataset can vary widely on the same benchmark without a clear explanation—we leave further investigation into this for future work.
@@ Appendix E Independent Benchmarks
In Figure 2, we present the performance of our best-performing routers on MT Bench as compared to Unify AI (UnifyAI, 2024) and Martian (Martian, 2024), two existing commercial offerings for LLM routing.
Here, we route between gpt-4-turbo-2024-04-09 (OpenAI, 2023) as $M_{s}$, and either mixtral-8x7b-instruct-v0.1 (Jiang et al., 2024) or llama-2-70b-chat (Touvron et al., 2023) as $M_{w}$ depending on which model each system supports. For Unify AI, we select the best-performing router configuration on the user dashboard and use it for benchmarking. For Martian, we optimize for performance and specify the maximum cost per million tokens as \$10.45, approximating this value using public inference costs (OpenAI, 2024; Together.AI, 2024) based on a 1:1 input:output token ratio so that 50% of calls are routed to GPT-4.
Both the matrix factorization router and causal LLM routers perform very competitively when trained on $\mathcal{D}_{\text{arena}}+\mathcal{D}_{\text{judge}}$, outperforming the commercial routing systems by achieving the same performance with up to 40% fewer calls routed to GPT-4.
@@ Appendix F Additional Plots
We include additional plots for the results presented in Section 5.1.